A boy has $84 and a girl has $12. How much does the boy need to give the girl in order for her to have the times the amount he has?
Thanks for the assist! I appreciate it!

Answer by ankor@dixie-net.com(22740)

A boy has $84 and a girl has $12.
How much does the boy need to give the girl in order for her to have the times the amount he has?
:
let a = amt of $ to make this happen
12 + a = 84 - a
a + a = 84 - 12
2a = 72
a = 72/2
a = $36
:
:
Confirm tjis
12 + 36 = $48
84 - 36 = $48